a meter scale is kept in a satellite moving with velocity 0.6w.r.t earth. calculate the length of the rod as measured by the observer from earth
1
Expert's answer
2020-12-14T12:12:53-0500

"\text{Lorentz factor according to}"

"l = l_0\sqrt{1-\frac{v^2}{c^2}}"

"v = 0.6c\text{ where } c\text{ speed of light }"

"l = l_0\sqrt{1-\frac{0.36c^2}{c^2}}= l_0*0.8 = 0.8m"

Answer:0.8m